OpenStudy (anonymous):

What are the approximate solutions of 5x2 − 7x = 1 rounded to the nearest hundredth?

OpenStudy (helder_edwin):

your equation can be rewritten into \[ \large 5x^2-7x-1=0 \] use the general formula. do u know it?

OpenStudy (anonymous):

well it's x=0.16 or x=1.24

OpenStudy (anonymous):

try using the quadratic formula yourself and checking with my answers.

OpenStudy (anonymous):

No real solutions x ≈ −0.65 and x ≈ 7.65 x ≈ −1.53 and x ≈ 0.13 x ≈ −0.13 and x ≈ 1.53 these are the answer choices.

OpenStudy (helder_edwin):

it is the 3rd one

OpenStudy (anonymous):

here a good site http://www.math.com/students/calculators/source/quadratic.htm

OpenStudy (helder_edwin):

but do u know how to get this

OpenStudy (anonymous):

just plug in

OpenStudy (anonymous):

no, could you please explain?

OpenStudy (helder_edwin):

the cuadratic formula gives u \[ \large x=\frac{7\pm\sqrt{49-4\cdot5\cdot(-1)}}{10}=\frac{7\pm\sqrt{69}}{10} \]

OpenStudy (anonymous):

okay thank you

OpenStudy (helder_edwin):

finally \[ \large \sqrt{69}\approx8.3066 \]